OpenStudy (anonymous):

Medal Give F (x) = x^2 + 3x -4 F (-6)

OpenStudy (usukidoll):

just let x = -6 and simplify the equation.

OpenStudy (anonymous):

can you show me

OpenStudy (usukidoll):

yeah sorry I crashed anyway f(x) -> f(-6) x = -6 so when we plug that into our equation we get F (x) = x^2 + 3x -4 F (-6) = (-6)^2 + 3(-6) -4 f(-6) = 36-18-4 f(-6) = 14

OpenStudy (anonymous):

thank you soo much :)

OpenStudy (usukidoll):

I caught someone on here asking for help on an online exam. It's against Openstudy's rules to do that. hahah 7 people viewing.

OpenStudy (anonymous):

oh wow this is very wrong